OpenStudy (anonymous):

sec^2(11pi/6) please help!!!!!!! evaluate

OpenStudy (anonymous):

we knowed that secx=1/cosx and the value of cos(11pi/6) = cos330 = sqrt(3)/2

OpenStudy (anonymous):

so, sec^2(11pi/6) = 1/{(sqrt(3)/2} = 1/(3/4) = 4/3